OpenStudy (anonymous):

Find the point P where the line x = 1 + t, y = 2t, z = -3t intersects the plane x + y - z = -1. Help please.

OpenStudy (anonymous):

I have -2t=-7+3s & 1+2t=5s & 3t=1+5s. Where do I go from here?

OpenStudy (anonymous):

If I isolate each variable, I get that t = 7/2 and s = 19/10, I am lost after that. I don't know what to do. Help please.

myininaya (myininaya):

hey

myininaya (myininaya):

did you try replace putting x+y-z=-1 in terms of t

myininaya (myininaya):

\[x+y-z=-1 \\ (1+t)+(2t)-(-3t)=-1\] Solve this for t

myininaya (myininaya):

then you can find (x,y,z) the point of intersection

myininaya (myininaya):

(x,y,z)=(1+t,2t,-3t)

myininaya (myininaya):

Are you here?

OpenStudy (anonymous):

yes I am

OpenStudy (anonymous):

so you are saying that the result would be (0, -2, 3). Is that right?

myininaya (myininaya):

what did you get for t?

OpenStudy (anonymous):

-1/3 = t

myininaya (myininaya):

ok isn't 1+(-1/3)=2/3

myininaya (myininaya):

that would be the x-coordinate

myininaya (myininaya):

and y=2t=2(-1/3)=-2/3 and z=-3t

myininaya (myininaya):

so z=?

OpenStudy (anonymous):

1

myininaya (myininaya):

yep

OpenStudy (anonymous):

okay I got it now. Thank you very much.

myininaya (myininaya):

np